# On the volume of K-semistable Fano manifolds

### v1

@@ 4. Volume estimates and minimal rational curves
$=$ $\sum_{m=0}^{\lfloor y/\ell\rfloor}\quad\sum_{\begin{subarray}{c}|I|+|J|=y-(\ell-1)\lfloor y/\ell\rfloor+m(\ell-1)\\ |I|=y-\ell\lfloor y/\ell\rfloor+m\ell\\ |J|=\lfloor y/\ell\rfloor-m\end{subarray}}a_{IJ}z^{\prime I}z^{\prime\prime J}.$
Assume that $W$ is another Zariski open set such that $W\cap Z$ is non-empty and has at most one nodal point. In particular, there are at most two irreducible components of $W\cap Z$. Fix one irreducible component $Z_{1}$ which is smooth, we can choose coordinates (after possibly shrinking $W$) such that
  1. (1)
    $W\cap Z_{1}=\{w_{1}=\cdots=w_{n-1}=0\}$
  2. (2)
    For any $21\leq i\leq d-$ 2, ${\rm d}w_{i}$ is a local generator of the $i$-th summand of ${\mathcal{O}}(-1)^{\oplus(d-2)}$ over $W\cap Z_{1}$.
  3. (3)
    For any $d-1\leq j\leq n-1$, ${\rm d}w_{j}$ is a local generator of the $j$-th summand of ${\mathcal{O}}^{\oplus(n-d+1)}$ over $W\cap Z_{1}$.
Near $U\cap W\cap Z$, the transition functions from $\mathbf{z}$ and $\mathbf{w}$ must be of the form:
$z_{k}=a_{k}(w_{n})w_{k}+R_{2},\quad 1\leq k\leq n-1$
where $a_{k}(w_{n})$ is a transition function of ${\mathcal{O}}_{{\mathbb{P}}^{1}}(-1)$ for $21\leq k\leq d-$ 2 and of ${\mathcal{O}}_{{\mathbb{P}}^{1}}$ for $d-1\leq k\leq n-1$, and $R_{2}$ consists of terms of degree at least 2 in $w_{1},\dots,w_{n-1}$. So near $W\cap Z_{1}$, we have $s=\sum_{|I|+|J|=y-(\ell-1)\lfloor y/\ell\rfloor}b_{IJ}(w_{n})w^{\prime I}w^{\prime\prime J}+R$ where $R$ are terms of (unweighted) total degree in $(w^{\prime},w^{\prime\prime})$ strictly greater than $y-(\ell-1)\lfloor y/\ell\rfloor$. We can deal with the other possible component of $W\cap Z$ using the same argument. So we see that $s$ induces a well-defined section in $H^{0}({\mathbb{P}}^{1},L^{k}\otimes{\rm Sym}^{y-\ell\lfloor y/\ell\rfloor}({\mathcal{O}}(-1)^{\oplus(d-2)})\otimes{\rm Sym}^{\lfloor y/\ell\rfloor}({\mathcal{O}}^{\oplus(n-d+1)}))$. This defines a linear map of vector spaces:
$H^{0}(X,L^{k}\otimes{\mathcal{I}}_{y}/{\mathcal{I}}_{y+1})\rightarrow H^{0}({\mathbb{P}}^{1},L^{k}\otimes{\rm Sym}^{y-\ell\lfloor y/\ell\rfloor}({\mathcal{O}}(-1)^{\oplus(d-2)})\otimes{\rm Sym}^{\lfloor y/\ell\rfloor}({\mathcal{O}}^{\oplus(n-d+1)})),$
whose kernel consists of elements $s\in H^{0}(X,L^{k}\otimes{\mathcal{I}}_{y}/{\mathcal{I}}_{y+1})$ that has expansion ($m\geq 1$ now):
@@
$s=\sum_{m=1}^{\lfloor y/\ell\rfloor}\quad\sum_{\begin{subarray}{c}|I|+|J|=y-(\ell-1)\lfloor y/\ell\rfloor+m(\ell-1)\\ |I|=y-\ell\lfloor y/\ell\rfloor+m\ell\\ |J|=\lfloor y/\ell\rfloor-m\end{subarray}}a_{IJ}z^{\prime I}z^{\prime\prime J},$
which by the same argument as above induces a section in $H^{0}({\mathbb{P}}^{1},L^{k}\otimes{\rm Sym}^{y-\ell\lfloor y/\ell\rfloor+\ell}({\mathcal{O}}(-1)^{\oplus(d-2)})\otimes{\rm Sym}^{\lfloor y/\ell\rfloor-1}({\mathcal{O}}^{\oplus(n-d+1)}))$. So we inductively get the estimate:
@@
$h^{0}(X,L^{k}\otimes{\mathcal{I}}_{y}/{\mathcal{I}}_{y+1})$$\leq$ $\sum_{m=0}^{\lfloor y/\ell\rfloor}h^{0}({\mathbb{P}}^{1},{\mathcal{O}}(kd)\otimes{\rm Sym}^{y-\ell\lfloor y/\ell\rfloor+m\ell}({\mathcal{O}}(-1)^{\oplus(d-2)})\otimes{\rm Sym}^{\lfloor y/\ell\rfloor-m}({\mathcal{O}}^{\oplus(n-d+1)}))$
@@
$\leq$ $\sum_{m=0}^{\lfloor y/\ell\rfloor}h^{0}({\mathbb{P}}^{1},{\mathcal{O}}(kd)\otimes{\rm Sym}^{\ell m}({\mathcal{O}}(-1)^{\oplus(d-2)})\otimes{\rm Sym}^{\lfloor y/\ell\rfloor-m}({\mathcal{O}}^{\oplus(n-d+1)})).$
@@
For the last inequality, we used the facts that $y-\ell\lfloor y/\ell\rfloor\geq 0$ and ${\mathcal{O}}(-1)$ is negative. Then the right-hand-side of (18) is bounded above by
@@
We will calculate the leading coefficient $b_{1}={\rm lim}_{k\rightarrow+\infty}I_{k}(x)/k^{n}$. First, we assume that $x\leq d$. Then we can expand:
@@
Similar calculation as above shows that this sum equals $b_{2}k^{n}/n!+O(k^{n-1})$ with $b_{2}$ equal to $C\cdot\mathbf{I}$ where $C=\frac{n!}{(d-3)!(n-d)!}\ell^{-(n-d+1)}$ and $\mathbf{I}$ is equal to:
@@
[…]
Set $T={\rm sup}\{x:V-\phi(x)=0\}$. If $T\leq d$, then we get $V=\phi(T)\leq\phi(d)=\ell^{-(n-d+1)}d^{n}(n-d+2)\leq d^{n}(n-d+2)$. Note that the number $d^{n}(n-d+2)$ is nothing but the volume of the Fano hypersurface in ${\mathbb{P}}^{n+1}$ of degree $n-d+2$ which is always less than $2n^{n}$ if $22<d<n+$ 2.
So we can assume $T\geq d$. The equation $\Phi(T)=(T-A)\phi(T)$ from (4) with $A=A_{X}(E)=(d-2)+(n-d+1)\ell$ becomes:
$\frac{1}{n-d+2}\int_{0}^{d}z^{d-3}(d-z)(T-z)^{n-d+2}\,{\rm d}z=(T-(d-2)-(n-d+1)\ell)\int_{0}^{d}z^{d-3}(d-z)(T-z)^{n-d+1}\,{\rm d}z.$
Even though the explicit solution of $T$ is not available, as $\ell\rightarrow+\infty$, the above equation can be written as:
$\frac{1}{n-d+2}T^{n-d+2}(1+O(T^{-1}))=(T-(n-d+1)\ell+O(1))T^{n-d+1}(1+O(T^{-1})).$
So we get the asymptotic expression for its solution: $T=T(\ell)=(n-d+2)\ell+O(1)$. So we can calculate that, up to $O(\ell^{-1})$, $\phi(T)$ is given by:
[…]
$=$ $\ell^{-(n-d+1)}\frac{n!}{(d-3)!(n-d+1)!}\left(\int_{0}^{1}t^{d-3}(1-t){\rm d}t\right)\cdot d^{d-1}\ell^{n-d+1}(n-d+2)^{n-d+1}$
$=$ $(n-d+2)^{n-d+1}\frac{n!}{(d-1)!(n-d+1)!}d^{d-1}={\rm vol}({\mathbb{P}}^{n-d+1}\times{\mathbb{P}}^{d-1}).$
Letting $\ell\rightarrow+\infty$, by Corollary 2.5, we have thus proved ${\rm vol}(X)=(-K_{X})^{n}\leq{\rm vol}({\mathbb{P}}^{d-1}\times{\mathbb{P}}^{n-d+1})$ and complete the proof of Theorem 1.4. ∎
@@
By using binomial expansion, one can get the following combinatorial expressions for $x\geq d$:
$\phi(x)$ $=$ $\ell^{-(n-d+1)}\sum_{j=d-1}^{n}\binom{n}{j}\cdot(j-d+2)\cdot d^{j}\cdot(x-d)^{n-j},$
$\Phi(x)$ $=$ $\ell^{-(n-d+1)}\sum_{j=d-1}^{n+1}\binom{n+1}{j}\frac{j-d+2}{n+1}d^{j}\cdot(x-d)^{n+1-j}.$
@@
When $d=n+1$, the above expression becomes: $\phi(x)=(n+1)^{n}$ (for $x\geq d$) which implies the volume estimate ${\rm vol}(X)\leq(n+1)^{n}={\rm vol}({\mathbb{P}}^{n})$.
When $d=n$, the above expressions become:
$\phi(x)$ $=$ $\ell^{-1}\left(n\cdot n^{n-1}\cdot(x-n)+2n^{n}\right)=\ell^{-1}n^{n}(x+2-n),$
$\Phi(x)$ $=$ $\frac{\ell^{-1}}{n+1}\left(\binom{n+1}{2}n^{n-1}(x-n)^{2}+(n+1)2n^{n}(x-n)+3n^{n+1}\right)$
$=$ $\frac{\ell^{-1}n^{n}}{2}\left(x^{2}-2(n-2)x+\frac{n}{(n+1)}(n-1)(n-2)\right).$
The equation $(T-(n-2)-\ell)\phi(T)=\Phi(T)$ is asymptotically given by:
$(T-\ell+O(1))T(1+O(T^{-1}))=\frac{1}{2}T^{2}(1+O(T^{-1}))$
which implies $T=2\ell+O(1)$. So $\phi(T)=\ell^{-1}n^{n}(2\ell+O(1)+2-n)\longrightarrow 2n^{n}$ as $\ell\rightarrow+\infty$ and we get an estimate: ${\rm vol}(X)\leq 2n^{n}={\rm vol}({\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1})={\rm vol}(Q^{n})$.

### v2

@@ 4.3. Minimal rational curves with degree $d\geq 3$
$\frac{s_{t}}{\sigma_{t}}$ $=$ $\sum_{|I|+2|J|=y}b^{t}_{IJ}(z_{n}){\mathbf{z}}^{\prime I}{\mathbf{z}}^{\prime\prime J}=\sum_{m=0}^{\lfloor y/2\rfloor}\sum_{\begin{subarray}{c}|I|+|J|=y-\lfloor y/2\rfloor+m\\ |I|=y-2\lfloor y/2\rfloor+2m\\ |J|=\lfloor y/2\rfloor-m\end{subarray}}b_{IJ}^{t}(z_{n}){\mathbf{z}}^{\prime I}{\mathbf{z}}^{\prime\prime J}$
which we call the initial term representation of $s$ over $U_{t}$. We set
@@
$\gamma_{t}([s]):=\sigma_{t}\cdot\sum_{\begin{subarray}{c}|I|=y-2\lfloor y/2\rfloor\\ |J|=\lfloor y/2\rfloor\\ |I|+|J|=y-\lfloor y/2\rfloor\end{subarray}}b^{t}_{IJ}(z_{n})({\rm d}{\mathbf{z}}^{\prime})^{\otimes I}({\rm d}{\mathbf{z}}^{\prime\prime})^{\otimes J}$
which defines a local section of $L^{k}\otimes{\rm Sym}^{y-2\lfloor y/2\rfloor}({\mathcal{O}}(-1)^{\oplus(d-2)})\otimes{\rm Sym}^{\lfloor y/2\rfloor}({\mathcal{O}}^{\oplus(n-d+1)}))$ over $B_{t}$. We claim that this section is well-defined over the whole ${\mathbb{P}}^{1}$, i.e. it is independent of $t\in{\mathbb{P}}^{1}$. For any other $t^{\prime}\in{\mathbb{P}}^{1}$, choose a sequence $\{t_{i};i=0,\dots,N\}$ and $B_{t_{i}}(\epsilon^{\prime})$ as before. We use the transition function (15) to get
$\frac{s_{t}}{\sigma_{t}}$ $=$ $\sum_{|I|+2|J|=y}b^{t}_{IJ}(g_{n}(\hat{z}_{n})+R_{1})(\mathbf{a}^{\prime}(z_{n})\hat{{\mathbf{z}}}^{\prime}+R_{2})^{I}(\mathbf{a}^{\prime\prime}(z_{n})\hat{{\mathbf{z}}}^{\prime\prime}+R_{2})^{J}$
$=$ $\sum_{|I|+2|J|=y}b^{t}_{IJ}(g_{n}(\hat{z}_{n}))(\mathbf{a}^{\prime}(z_{n})\hat{{\mathbf{z}}}^{\prime})^{I}(\mathbf{a}^{\prime\prime}(z_{n})\hat{{\mathbf{z}}}^{\prime\prime})^{J}+\sum_{|I|+2|J|=y}b^{t}_{IJ}(g_{n}(\hat{z}_{n}))(\mathbf{a}^{\prime}(z_{n})\hat{{\mathbf{z}}})^{\prime I}R^{\prime J}_{2}+R^{w}_{y+1}$
$=:$ $\frac{s_{t^{\prime}}}{\sigma_{t^{\prime}}}\frac{\sigma_{t^{\prime}}}{\sigma_{t}}+\tilde{R}^{w}_{y+1}$
where $R^{\prime}_{2}$ is quadratic in $\hat{z}_{1},\dots,\hat{z}_{d-1}$, $R^{w}_{y+1}$ consists of terms of weighted degree $2|I|+J\geq y+1$ and $\tilde{R}^{w}_{y+1}$ consists of terms with $2|I|+J\geq y$ and $|I|+|J|>y-2\lfloor y/2\rfloor$. As a consequence, we get:
$\gamma_{t^{\prime}}([s])$ $=$ $\sigma_{t}\cdot\sum_{\begin{subarray}{c}|I|=y-2\lfloor y/2\rfloor\\ |J|=\lfloor y/2\rfloor\end{subarray}}b^{t}_{IJ}(g_{n}(\hat{z}_{n}))(a^{\prime}(z_{n}){\rm d}\hat{{\mathbf{z}}}^{\prime})^{\otimes I}(a^{\prime\prime}(z_{n})\hat{{\mathbf{z}}}^{\prime\prime})^{\otimes J}$
which coincides with $\gamma_{t}([s])$ on $B_{t}\cap B_{t^{\prime}}$.
So we get a linear map of vector spaces:
@@
$\gamma:H^{0}(X,L^{k}\otimes{\mathcal{I}}_{y}/{\mathcal{I}}_{y+1})\rightarrow H^{0}({\mathbb{P}}^{1},L^{k}\otimes{\rm Sym}^{y-2\lfloor y/2\rfloor}({\mathcal{O}}(-1)^{\oplus(d-2)})\otimes{\rm Sym}^{\lfloor y/2\rfloor}({\mathcal{O}}^{\oplus(n-d+1)})),$
whose kernel consists of elements $s]\in H^{0}(X,L^{k}\otimes{\mathcal{I}}_{y}/{\mathcal{I}}_{y+1})$ that has the initial term representation given by ($m\geq 1$ now):
$\frac{s_{t}}{\sigma_{t}}=\sum_{m=1}^{\lfloor y/2\rfloor}\quad\sum_{\begin{subarray}{c}|I|+|J|=y-\lfloor y/2\rfloor+m\\ |I|=y-2\lfloor y/2\rfloor+2m\\ |J|=\lfloor y/2\rfloor-m\end{subarray}}b_{IJ}(z_{n}){\mathbf{z}}^{\prime I}{\mathbf{z}}^{\prime\prime J},$
@@
which by the same argument as above induces a section in $H^{0}({\mathbb{P}}^{1},L^{k}\otimes{\rm Sym}^{y-2\lfloor y/2\rfloor+2}({\mathcal{O}}(-1)^{\oplus(d-2)})\otimes{\rm Sym}^{\lfloor y/2\rfloor-1}({\mathcal{O}}^{\oplus(n-d+1)}))$. So we inductively get the estimate:
$h^{0}(X,L^{k}\otimes{\mathcal{I}}_{y}/{\mathcal{I}}_{y+1})$
$\leq$ $\sum_{m=0}^{\lfloor y/2\rfloor}h^{0}({\mathbb{P}}^{1},{\mathcal{O}}(kd)\otimes{\rm Sym}^{y-2\lfloor y/2\rfloor+2m}({\mathcal{O}}(-1)^{\oplus(d-2)})\otimes{\rm Sym}^{\lfloor y/2\rfloor-m}({\mathcal{O}}^{\oplus(n-d+1)}))$
(21) $\leq$ $\sum_{m=0}^{\lfloor y/2\rfloor}h^{0}({\mathbb{P}}^{1},{\mathcal{O}}(kd)\otimes{\rm Sym}^{2m}({\mathcal{O}}(-1)^{\oplus(d-2)})\otimes{\rm Sym}^{\lfloor y/2\rfloor-m}({\mathcal{O}}^{\oplus(n-d+1)})).$
@@
For the last inequality, we used the facts that $y-2\lfloor y/2\rfloor\geq 0$ and ${\mathcal{O}}(-1)$ is negative.
Remark 4.4.
In the first version of this preprint, we defined valuations $v_{\ell}$ for any $\ell\geq 1\in{\mathbb{N}}$ and claimed that the above estimate is true for any $\ell\geq 1\in{\mathbb{N}}$. Ziquan Zhuang pointed out that for $\ell\geq 3$, this is not true. The reason is that while $v_{1}$ and $v_{2}$ are well-defined globally, the valuation $v_{\ell}$ for $\ell\geq 3$ depends on a collection of coordinate charts that only cover a dense open subset $f({\mathbb{P}}^{1})$ (the complement of finite many points), which causes the failure of the above estimate. For $v_{1}$ and $v_{2}$ the coordinates charts we constructed cover the whole $f({\mathbb{P}}^{1})$ and there is no such an issue. By a similar argument in Subsection 4.1, a sufficient condition for $v_{\ell}$ to be defined globally is that the transition functions between coordinate charts covering the whole $f({\mathbb{P}}^{1})$ are of the form:
$z_{k}=a_{k}(\hat{z}_{n})\hat{z}_{k}+R_{\ell},1\leq k\leq n-1;z_{n}=g_{n}(\hat{z}_{n})+R_{\ell-1}$
where $R_{\ell}$ (resp. $R_{\ell-1}$) consists of terms of degree at least $\ell$ (resp. at least $\ell-1$) in $\hat{z}_{1},\dots,\hat{z}_{n-1}$. When the minimal rational curve is embedded, this condition is satisfied exactly when the rational curve is $(\ell-2)$-comfortably embedded in the sense of Abate-Bracci-Tovena (see [ABT09, Remark 3.3]). The latter condition is a refinement and a slight weakening of Grauert’s $(\ell-1)$-linearizability condition. From this point of view, the valuation $v_{1}$, $v_{2}$ are always globally defined since any minimal rational curve is always $1$-linearizable, and in particular (trivially) 0-comfortably embedded.
By the similar (and simpler) argument as above, we get the similar estimate for the valuation $v_{1}$ (i.e $v_{\ell}$ from section 4.1 with $\ell=1$):
$h^{0}(X,L^{k}\otimes{\mathcal{I}}_{y}(v_{1})/{\mathcal{I}}_{y+1}(v_{1}))$ $\leq$ $\sum_{m=0}^{y}h^{0}({\mathbb{P}}^{1},{\mathcal{O}}(kd)\otimes{\rm Sym}^{m}({\mathcal{O}}(-1)^{\oplus(d-2)})\otimes{\rm Sym}^{y-m}({\mathcal{O}}^{\oplus(n-d+1)})).$
By (16) and (21), for each $v_{\ell}$ with $\ell\in\{1,2\}$, we get that $h^{0}(X,L^{k}\otimes{\mathcal{O}}_{X}/{\mathcal{I}}_{xk}(v_{\ell}))$ is bounded from above by:  [^1] Our main interest will be the case $\ell=2$. We include the parameter $\ell=1$ for comparison later.
@@
We will calculate the leading coefficient $b_{1}={\rm lim}_{k\rightarrow+\infty}(I_{k}(x)\cdot n!)/k^{n}$. First, we assume that $x\leq d$. Then we can expand:
@@
Similar calculation as above shows that this sum equals $b^{\prime}_{1}k^{n}/n!+O(k^{n-1})$ with $b^{\prime}_{1}$ equal to $C\cdot\mathbf{I}$ where $C=\frac{n!}{(d-3)!(n-d)!}\ell^{-(n-d+1)}$ and $\mathbf{I}$ is equal to:
@@
(26) $\Phi(x)=\int_{0}^{x}\phi(t)dt=\left\{\begin{array}[]{ll}\frac{\ell^{-(n-d+1)}x^{n}}{n+1}(d(n+1)-(d-2)x)&0\leq x\leq d\\ \frac{\ell^{-(n-d+1)}n!}{(d-3)!(n-d+2)!}\int_{0}^{\rm d}z^{d-3}(d-z)(x-z)^{n-d+2}{\rm d}z&x\geq d.\end{array}\right$.
By using the binomial expansion:
$(x-z)^{n-d+1}=(x-d+d-z)^{n-d+1}=\sum_{j=0}^{n-d+1}\binom{n-d+1}{j}(d-z)^{j}(x-d)^{n-1+d-j}.$
it is straightforward to get the following combinatorial expressions when $x\geq d$:
(27) $\phi(x)$ $=$ $\ell^{-(n-d+1)}\sum_{j=0}^{n-d+1}\binom{n}{j}\cdot(n-d+2-j)d^{n-j}(x-d)^{j},$
(28) $\Phi(x)$ $=$ $\ell^{-(n-d+1)}\sum_{j=0}^{n-d+2}\binom{n+1}{j}\frac{n-d+3-j}{n+1}d^{n+1-j}(x-d)^{j}$
Set $T={\rm sup}\{x:V-\phi(x)=0\}$. If $T\leq d$, then we get $V=\phi(T)\leq\phi(d)=\ell^{-(n-d+1)}d^{n}(n-d+2)\leq d^{n}(n-d+2)$. Note that the number $d^{n}(n-d+2)$ is nothing but the volume of the Fano hypersurface in ${\mathbb{P}}^{n+1}$ of degree nd+2n-d+2 which is always less than $2n^{n}$ if $2<d<n+2$.
Remark 4.5.
Setting $d=2$ in the formula (27)-(28), we also recover functions from Example 2.7 (the $\ell^{1-n}$ factor comes the fact that here we are using the $\ell^{\oplus(n-1)}$ blowup):
$\phi(x)=2n\cdot\ell^{1-n}x^{n-1},\quad\Phi(x)=2\ell^{1-n}x^{n}.$
So we can assume $T\geq d$. Now we set $\ell=2$ and $A=A_{X}(v_{2})=(d-2)+2(n-d+1)=2n-d$. The equation $\Phi(T)=(T-A)\phi(T)$ from (4) becomes:
(29) $\frac{1}{n-d+2}\int_{0}^{d}z^{d-3}(d-z)(T-z)^{n-d+2}\,{\rm d}z=(T-(2n-d))\int_{0}^{d}z^{d-3}(d-z)(T-z)^{n-d+1}\,{\rm d}z.$
By Corollary 2.5, there exists a unique solution $T=T(d,n)$ to the above equation.
@@
When $d=n+1$, the expression (23) becomes: $\phi(x)=(n+1)nx^{n-1}-(n-1)x^{n}$ if $x\leq n+1$ and $\phi(x)=(n+1)^{n}$ (for $x\geq n+1$). This implies the (un-interesting) volume estimate ${\rm vol}(X)\leq(n+1)^{n}={\rm vol}({\mathbb{P}}^{n})$.
Example 4.7.
When $d=n$, the expressions (27)-(28) become:
(30) $\phi(x)$ $=$ $\ell^{-1}\left(n\cdot n^{n-1}\cdot(x-n)+2n^{n}\right)=\ell^{-1}n^{n}(x+2-n),$
$\Phi(x)$ $=$ $\frac{\ell^{-1}n^{n}}{2}\left(x^{2}-2(n-2)x+\frac{n}{(n+1)}(n-1)(n-2)\right).$
When $d=n$ and for the case of $(1^{\oplus(d-2)},\ell^{\oplus(n-d+1)})$-weighted blowup, equation $(T-A(v_{\ell}))\phi(T)=\Phi(T)$ is equivalent to $T^{2}-2(n-2+\ell)T+\left(\frac{n(n-2)(n+3)}{2(n+1)}+(2\ell-4)\right)=0$. One can solve
$T=\left\{\begin{array}[]{ll}(n-1)+\sqrt{\frac{3(n-1)}{n+1}}&\text{for }\ell=1;\\ n+\sqrt{\frac{6n}{n+1}}&\text{for }\ell=2.\end{array}\right$.
So we get the volume upper bound ${\rm vol}(X)\leq\phi_{\ell}(T)$ which is given by:
$\phi_{\ell}(T)=\left\{\begin{array}[]{ll}n^{n}(T-n+2)=n^{n}(1+\sqrt{\frac{3(n-1)}{n+1}})&\text{for }\ell=1\\ \frac{1}{2}n^{n}(T-n+2)=n^{n}\left(1+\sqrt{\frac{3n}{2(n+1)}}\right)&\text{for }\ell=2.\end{array}\right$.
Note that when $n=2$, $\phi_{1}(T)=\phi_{2}(T)={\rm vol}(Q)=8$. When $n\geq 3$, $\phi_{1}(T)>\phi_{2}(T)>2n^{n}={\rm vol}(Q)$. To get the sharp upper bound of ${\rm vol}(X)=(-K_{X})^{n}$ for all $d=n\geq 3$, we are going to use the classification results stated in Theorem 2.9.
By Corollary 2.5, Proposition 4.3 is equivalent to the following proposition.
Proposition 4.8.
Let $3\leq d\leq n-1$ and $T$ be the solution to equation (29). We have $\phi(T)<2n^{n}$ where $\phi$ was defined in (23). As a consequence, we know that if $X$ is a K-semistable Fano manifold $3\leq l_{X}=d\leq n-2$, then ${\rm vol}(X)<2n^{n}$.
Remark 4.9.
Example 4.7 shows that for $d=n$ the estimate $\phi(T)<2n^{n}$ is not true. Surprisingly, Proposition 4.8 implies that the $(1^{\oplus(d-2)},2^{\oplus(n-d+1)})$ weighted blow-up is sufficient to establish the sharp estimate for all $3\leq d\leq n-1$. Even though the statement appears elementary, due to the complicated expansion from (27)-(28) it seems not so easy to find a simple proof for all pairs of positive integers $(d,n)$ satisfying $3\leq d\leq n-1$. Motivated by some numerical calculations, we managed to get the desired estimate using some tools from calculus and probability theory. See Figure 1 for the numerical illustration for $n=7$ and $2\leq d\leq 7$.
The rest of this subsection is devoted to explain the proof of Proposition 4.8. We first consider the case d=n1d=n-1. Note that equation $(T-A_{X}(E))\phi(T)=\Phi(T)$ appeared in Corollary 2.5 is a degree nd+2n-d+2 polynomial equation. When d=n1d=n-1, the corresponding equation is a cubic equation in $T$ for which we can explicitly compute the root $T$ using Cardano’s formula.
Proposition 4.10.
Proposition 4.8 holds when d=n1d=n-1.
Proof.
When d=n1d=n-1 and $\ell=2$, the expressions from (27)-(28) become:
$\phi(x)$ $=$ $\frac{1}{2^{2}}\left(\frac{n(n-1)^{n-1}}{2}(x-n+1)^{2}+2n(n-1)^{n-1}(x-n+1)+3(n-1)^{n}\right),$
$\Phi(x)$ $=$ $\frac{1}{2^{2}}(\frac{n(n-1)^{n-1}}{6}(x-n+1)^{3}+n(n-1)^{n-1}(x-n+1)^{2}$
$\,+3(n-1)^{n}(x-n+1)+\frac{4}{n+1}(n-1)^{n+1}).$
The equation $(T-A(E))\phi(T)=\Phi(T)$ is equivalent to $\tau^{3}-12\tau-\frac{6(n-1)(5n+1)}{n(n+1)}=0$ where $\tau=T-n+1$. By Cardano’s formula,
$\tau=\sqrt[3]{-\frac{q}{2}+\sqrt{\frac{q^{2}}{4}-64}}+\sqrt[3]{-\frac{q}{2}-\sqrt{\frac{q^{2}}{4}-64}},$
where $q=-\frac{6(n-1)(5n+1)}{n(n+1)}$. Hence, ${\rm vol}(X)\leq\phi(T)=\frac{n(n-1)^{n-1}}{8}((\tau+2)^{2}+2-\frac{6}{n})\eqqcolon F(n)$. We claim that for $n\geq 4$, we have $F(n)<2n^{n}={\rm vol}(Q)$.
When $n\geq 4$, note that the sequence $\tau(n)$ is monotone increasing with respect to $n$ and ${\rm lim}_{n\rightarrow+\infty}\tau(n)=\sqrt[3]{15+\sqrt{161}}+\sqrt[3]{15-\sqrt{161}}\eqqcolon\tau_{\infty}$. So, in particular, we have $\tau\leq\tau_{\infty}$. We further consider
$\frac{F(n)}{2n^{n}}$ $\leq$ $\frac{n(n-1)^{n-1}}{8\cdot 2n^{n}}\left((\tau_{\infty}+2)^{2}+2-\frac{6}{n}\right)$
$\leq$ $\frac{1}{16\left(1+\frac{1}{n-1}\right)^{n-1}}\left(\left(\sqrt[3]{15+\sqrt{161}}+\sqrt[3]{15-\sqrt{161}}+2\right)^{2}+2\right)\eqqcolon\gamma(n).$
Note that $\gamma(n)$ is monotone decreasing with respect to $n$ and ${\rm lim}_{n\rightarrow\infty}\gamma(n)=\frac{1}{16e}(\sqrt[3]{15+\sqrt{161}}+\sqrt[3]{15-\sqrt{161}}+2)^{2}<1$. So $\gamma(n)<1$ when $n$ is sufficiently large. By numerical method, we can show that $\gamma(n)<1$ when $n\geq 19$. For the case when $4\leq n\leq 18$, by direct computation, one can check that $F(n)<2n^{n}$. ∎
Remark 4.11.
If we use the valuation $v_{1}$ in the case of d=n1d=n-1, we can do a similar computation as in the proof of Proposition 4.10, then
$\tau=T-n+1=-1+\sqrt[3]{-1+\frac{q}{2}+\sqrt{\frac{q^{2}}{4}-q}}+\sqrt[3]{-1+\frac{q}{2}-\sqrt{\frac{q^{2}}{4}-q}}$
where $q=\frac{12(n-1)^{2}}{n(n+1)}$. In this case, $\phi(T)=\frac{n(n-1)^{n-1}}{2}\tau^{2}+2n(n-1)^{n-1}\tau+3(n-1)^{n}>2n^{n}={\rm vol}(Q)$ when $n\geq 4$. This shows that the use of $(1^{\oplus(d-2)},2^{\oplus(n-d+1)})$-weighted blowup is crucial for the proof of Proposition 4.3 and hence for Theorem 1.1.
Completion of proof of Proposition 4.8.
Because of Proposition 4.10, we just need to prove Proposition 4.8 when $3\leq d\leq n-2$. For simplicity, set
$\Psi(x)$ $=$ $(x-(2n-d))\phi(x)-\Phi(x).$
Note that $\Psi^{\prime}(x)=(x-(2n-d))\phi^{\prime}(x)$ since $\phi(x)=\Phi^{\prime}(x)$. Since $\phi^{\prime}(x)>0$ for $x>0$, we have the two useful facts:
  • $\Psi(x)$ is negative for $x\in(0,T)$ and positive for $x\in(T,+\infty)$, where $T$ is the unique solution to $\Psi(T)=0$ (see Corollary 2.5).
  • x>2nd+1x>2n-d+1, $\Psi^{\prime}(x)>\phi^{\prime}(x)$.
Now suppose there exists $x_{1}\geq 2n-d+1$ such that
$2n^{n}-\phi(x_{1})+\Psi(x_{1})>0,$
or equivalently $\phi(x_{1})<2n^{n}-\Psi(x_{1})$. If $\Psi(x_{1})\geq 0$, then $T\leq x_{1}$ and $\phi(T)\leq\phi(x_{1})<2n^{n}$. If $\Psi(x_{1})<0$, then $T>x_{1}$ and since $\Psi(T)=0$,
$\phi(T)$ $=$ $\phi(x_{1})+\int_{x_{1}}^{T}\phi^{\prime}(x){\rm d}x<\phi(x_{1})+\int_{x_{1}}^{T}\Psi^{\prime}(x){\rm d}x$
$=$ $\phi(x_{1})+\Psi(T)-\Psi(x_{1})<2n^{n}.$
By this simple calculus argument, we see that the proposition 4.8 is proved once we prove (31) for some $x_{1}\geq 2n-d+1$. Now the main claim is that (31) indeed holds when $3\leq d\leq n-2$ and $x_{1}=2n-d+2$. In other words, we need to show that for any $3\leq d\leq n-2$, the following inequality is true:
(31) $\phi(2n-d+2)-\Psi(2n-d+2)=\Phi(2n-d+2)-\phi(2n-d+2)<2n^{n}.$
Since the proof depends on long estimates, we put the proof in the Appendix A. ∎
4.4. Completion of the proof of Theorem 1.1